HELP PLEASE!!
What is the rule for the nth term :
d = -12 and a(subscript 1) = 80
Thanks very much
llo

Answer by stanbon(75887)

What is the rule for the nth term :
d = -12 and a(subscript 1) = 80
------------------
General Rule: a(n)= a(1)+(n-1)d
Your Problem:
a(n)=80+(n-1)*(-12)
a(n)=80-12n+12
a(n)=-12n+92
=====================
Cheers,
Stan H.